OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Gabriella went skiing. She paid $35 to rent skis and $15 an hour to ski. If she paid a total of $95, how many hours did she ski? Write an equation to represent the situation and solve it.

jagr2713 (jagr2713):

95-35=60/15=4